# Thermal Inertia Controls on Titan’s Surface Temperature and Planetary Boundary Layer Structure

### v1

@@ 3.2 Seasonal Variations
Figure 2 shows the annual zonal-mean surface temperatures as a function of latitude and time for the four simulation cases. As previously reported in MacKenzie et al. (2019)—which explored cases using a uniform thermal inertia of [TIU] (representing porous ice regolith; Tokano, 2005), a globally averaged value of [TIU], and a spatially varying thermal inertia based on the “convective lakes” scenario—seasonal surface temperatures in our Cases 1, 2, and 4 similarly exhibit little to no sensitivity to thermal inertia. In Case 4, the surface is dominated by low thermal inertia terrains: Plains make up approximately 65% of the surface area ( [TIU]), dunes 18% ( [TIU]), and labyrinth terrain 3% ( [TIU]). In contrast, high thermal inertia terrain—hummocky regions with [TIU]—comprise only 14% of the surface (see Figure 2 from MacKenzie et al., 2019). Among the simulation cases, Case 3 alone exhibits a distinct structure compared to the other three cases, indicating that seasonal surface temperature variations begin to show a noticeable sensitivity to thermal inertia only when the thermal inertia is sufficiently high globally.
@@
This behavior can be qualitatively understood using Eqs. (6) and (7). For an annual forcing frequency of [s-1], the corresponding values of for Cases 1, 2, and 3 are approximately 0.014, 0.045, and 0.14 [WmK-1], respectively. Among these, only Case 3 exceeds the atmospheric damping coefficient [WmK-1], obtained as the annual mean, indicating that subsurface heat conduction becomes significant only when thermal inertia is sufficiently high. On seasonal timescales, subsurface conduction is generally less important than atmospheric damping primarily due to Titan’s long orbital period (29.5 Earth years), which yields a small . This stands in contrast to the diurnal regime, wherein subsurface heat conduction can dominate the surface temperature variability at intermediate values of thermal inertia.
@@
Figure 3 presents the seasonal evolution of surface energy balance components and surface temperature near the Huygens landing site for the four simulation cases. Across all cases, the net SW flux varies seasonally from approximately 0.20 to 0.37 [Wm-2], increasing from winter through summer. The sensible heat flux exhibits a broadly similar seasonal pattern, ranging from 0.15 to 0.30 [Wm-2], but with lower thermal inertia cases showing larger fluctuation amplitudes. This enhanced fluctuation arises because, when subsurface heat conduction is inefficient due to low thermal inertia, a stronger sensible heat flux is required to remove the net energy at the surface, as also occurs at diurnal timescales. The net IR flux remains relatively steady at approximately 0.07 [Wm-2] across seasons and simulation cases. The resulting net energy flux is smaller in the lower thermal inertia cases, where a stronger sensible heat flux efficiently removes much of the incoming surface energy.
@@
Surface temperature variations across seasons remain modest—under 1 [K]—and exhibit broadly similar patterns across all simulation cases. However, the amplitude of these variations is more strongly damped for higher thermal inertia. Applying Eqs. (6) and (7) to the simulated seasonal net energy flux amplitudes of , , and [Wm-2] for Cases 1, 2, and 3, respectively, yields theoretical temperature variations of , 0.57, and 0.51 [K]. Here, is derived from means calculated over the final 30 simulation years where 10-Titan-day moving averages are used to remove short-term variations. These estimates are roughly consistent with the simulated surface temperature variations of , , and [K], though some discrepancies remain. These differences likely stem from the assumption of a constant across seasons, whereas in reality the horizontal wind speed exhibits seasonal variability in the simulations. Despite discrepancies between and due to the simplification, Eqs. (6) and (7) effectively capture the dominant role of atmospheric damping in regulating surface temperature variations at seasonal timescales. For instance, omitting the atmospheric damping term leads to theoretical predictions of , 1.22, and 0.70 [K], which significantly overestimate the temperature response compared to the simulations. For Case 4, the [Wm-2] yields a theoretical surface temperature variation of [K], nearly identical to that of Case 2.
@@ 4.1 Diurnal Variations
Figure 4 presents the diurnal evolution of the PBL height across the four simulation cases. The left panels in each case show the temporal evolution of potential temperature profiles from 9:00 AM to 3:00 PM local solar time, compared with in-situ measurements from the Huygens probe (Fulchignoni et al., 2005). The right panels display the corresponding temperature diffusion tendencies predicted by the model’s PBL parameterization. Consistent with the simulated surface temperature variations of [K] across the cases, temporal changes in the PBL structure are evident in both the potential temperature profiles and diffusion tendencies. The PBL depth increases from approximately 500 [m] in the late morning to about 1,000 [m] by mid-afternoon.
Temperature diffusion tendencies weaken with increasing thermal inertia, consistent with the reduction in sensible heat flux shown in Figure 1. Despite these differences, the simulated PBL heights remain similar across cases, with maximum depths reaching approximately 1,000 [m] regardless of thermal inertia. However, notable differences emerge in the upper troposphere—around 100 [mbar], corresponding to an altitude of approximately 50 [km]—where the simulations diverge. Specifically, Case 1, representing the low thermal inertia scenario, exhibits a higher tropopause temperature of about 71 [K], compared to 70 [K] in Case 3 (high thermal inertia), with Case 2 falling in between (not shown).
This difference reflects the cumulative effect of sensible heat flux over time: low thermal inertia surfaces promote greater upward energy transfer. As a result, the thermal influence extends throughout the troposphere, raising the equilibrium temperature near the tropopause. In contrast, high thermal inertia reduces sensible heat flux, limiting the energy available to warm the free atmosphere. In summary, while thermal structure differences among cases appear minimal at the diurnal timescale, they become more pronounced over longer periods due to the cumulative adjustment of the atmosphere to surface energy exchange. Relatedly, the influence of thermal inertia differences on seasonal wind structures is discussed in the following subsection.
@@ 4.2 Seasonal Variations
Figure 5 presents the seasonal evolution of the PBL height across the four simulation cases. The left panels in each case display global zonal means, while the right panels show seasonal variations near the Huygens landing site, alongside the temperature diffusion tendency. The PBL height is diagnosed as the altitude of the model layer at which the vertical diffusion coefficient first drops below a threshold of 0.01 [m2s-1], a criterion commonly employed within the Mellor-Yamada scheme to define the extent of the near-surface adiabatic layer. These diagnosed PBL heights reflect variations in surface heating and the resulting convective mixing, as corroborated by the simulated temperature diffusion tendencies in our simulations.
@@
The global seasonal structure of the PBL closely mirrors the surface temperature patterns shown in Figure 1, indicating that PBL depth variations are primarily governed by the seasonal migration of regions with peak surface heating. The hemispheric asymmetry in the PBL structure is evident, as it is in the surface temperature, due to orbital eccentricity. In Case 1, the diagnosed PBL height at the Huygens landing site varies from approximately 300 [m] during southern winter to about 1,000–1,500 [m] between the southern vernal () and autumnal () equinoxes.
Unlike at diurnal timescales, diffusion tendencies at seasonal timescales exhibit clear sensitivity to thermal inertia: maximum PBL heights reach approximately 2,000 [m] and 3,000 [m] in Cases 2 and 3, respectively. The elevated PBL depth for higher thermal inertia corresponds to a reduced contrast between the minimum and maximum surface temperature, as shown in Figure 3.
@@
Figure 6 presents global seasonal mean maps of the diagnosed PBL depth for four distinct seasons, based on Case 4. The maximum depth is around 1,800 [m], and structural differences across longitudes are apparent due to spatial variations in thermal inertia. For instance, an elevated PBL depth is observed in the Xanadu region (approximately 250°E near the equator) across the seasons, where the assumed thermal inertia corresponds to hummocky terrain (1,962 [TIU]), compared to dune fields (236 [TIU]) and plains (601 [TIU]) at similar latitudes. The imprint of this enhancement is especially pronounced during the southern summer and winter solstices. Additionally, elevated PBL depths are evident in polar regions—near 90°S, 180°E during southern summer and near 70°N, 40°E during southern winter—attributable to the presence of convective lakes with extremely high thermal inertia (18,090 [TIU]). These results suggest that locally varying thermal inertia strongly influences the seasonal evolution of the PBL depth.
@@
Figure 7 shows the zonal and meridional wind profiles at the model grid point closest to the Huygens landing site for the four simulation cases, compared with the Huygens probe observations (Bird et al., 2005; Karkoschka, 2016). The seasonal mean is computed as a 30 Titan-day average centered around midsummer. During the period, the simulated wind profiles exhibit minimal diurnal variation (not shown). Across all cases, the simulations show good agreement with the observed wind structure, capturing key features of both the zonal and meridional components. The simulated near-surface wind magnitudes are on the order of – [ms-1], consistent with results from previous Titan model intercomparison studies (Lora et al., 2019).
In the case of zonal winds, the upper atmosphere on Titan is dominated by eastward (superrotating) flow, driven by the upward and equatorward transport of angular momentum by the mean flow and eddies, respectively (Flasar & Achterberg, 2009; Lewis et al., 2023; Lombardo & Lora, 2023). Near the surface, frictional drag keeps the flow closer to Titan’s solid body rotation. The net gain of angular momentum from the surface at low latitudes necessitates mean westward (retrograde) winds near the surface, a feature consistently reproduced across all four simulations in our study and in prior Titan GCM studies (e.g., Lora et al., 2019). Although differences among simulation cases are minimal near the surface, substantial variations emerge in the stratosphere. In Cases 1, 2, and 4, maximum wind speeds reach approximately 150 [ms-1], whereas in Case 3 they peak at only 60 [ms-1] (not shown). This suggests that differences in global surface temperature structure—driven by varying thermal inertia—can influence wind patterns even at high altitudes. While this implies that surface-induced thermal structures may propagate upward and modulate stratospheric circulation, a detailed analysis of the vertical momentum transfer and dynamical response is beyond the scope of this study.
@@
For meridional winds, the flow near the surface is predominantly southward, transitioning to northward at altitudes around 2 [km]. This vertical wind reversal suggests the presence of a cross-equatorial Hadley circulation, characterized by rising motion in the southern summer hemisphere and descending motion in the northern winter hemisphere. As a result, there is northward return flow aloft and a compensating southward return flow near the surface at the latitude of the Huygens landing site (Lora et al., 2025). Wind magnitudes—both zonal and meridional—are generally smaller in the higher thermal inertia cases, although the differences are minimal. This likely reflects weaker temperature gradients, as higher thermal inertia surfaces dampen surface temperature contrasts (Figure 2), thereby reducing the thermal forcing that drives the atmospheric circulation.
@@ 5 PBL Characteristics at the Dragonfly Landing Site
In this section, we present the same analysis of surface temperature and PBL structure as in the previous sections, but for the planned Dragonfly landing site near Selk Crater (E, N) during southern midsummer () (Barnes et al., 2021; Lorenz et al., 2021). For Case 4, the dunes terrain type, with a thermal inertia of [TIU], is assumed at the surface (MacKenzie et al., 2019).
Figure 8 shows the diurnal energy balance components at the Dragonfly landing site. Maximum insolation is reduced compared to the Huygens landing site, as the Dragonfly site experiences northern winter conditions while the Huygens site corresponds to southern summer. This results in a lower mean surface temperature at the Dragonfly site, reduced by approximately 0.1 [K] relative to the Huygens site. The amplitude of diurnal surface temperature variations remains comparable to that at the Huygens site. For Case 4, the simulated diurnal amplitude is [K].
Figure 9 shows the diurnal PBL structures at the Dragonfly landing site. Due to the reduced surface energy input and lower temperatures during this winter season—compared to the midsummer conditions at the Huygens landing site—the diagnosed PBL depths are shallower, reaching a maximum of approximately 500 [m], in contrast to 1,000 [m] at the Huygens site. Nevertheless, the overall temperature structures remain similar to those observed at the Huygens landing site.
Figure 10 shows the seasonal evolution of PBL depth at the Dragonfly landing site. As with the Huygens landing site, higher thermal inertia surfaces lead to elevated PBL depths. However, a key difference is observed in the seasonal symmetry. At the Dragonfly site, the PBL depth exhibits a more symmetric seasonal pattern: maxima occur near the equinoxes ( and ), and minima near the solstices ( and ). In contrast, the Huygens site displays an asymmetric pattern: a pronounced minimum at the southern winter solstice (), with a higher but still suppressed depth at the southern summer solstice (), relative to the equinoctial maxima. This asymmetry is attributed to Saturn’s orbital eccentricity, as illustrated in Figure 2 and the left panels of Figure 5.

### v2

@@ 3.2 Seasonal Variations
Figure 2 shows the annual zonal-mean surface temperatures as a function of latitude and time for the four simulation cases. As previously reported in MacKenzie et al. (2019)—which explored cases using a uniform thermal inertia of TIU (representing porous ice regolith; Tokano, 2005), a globally averaged value of TIU, and a spatially varying thermal inertia based on the “convective lakes” scenario—seasonal surface temperatures in our Cases 1, 2, and 4 similarly exhibit little to no sensitivity to thermal inertia. In Case 4, the surface is dominated by low thermal inertia terrains: Plains make up approximately 65% of the surface area ( TIU), dunes 18% ( TIU), and labyrinth terrain 3% ( TIU). In contrast, high thermal inertia terrain—hummocky regions with TIU—comprise only 14% of the surface (see Figure 2 from MacKenzie et al., 2019). Among our simulation cases, Case 3 alone exhibits a distinct structure compared to the other three, indicating that seasonal surface temperature variations begin to show a noticeable sensitivity to thermal inertia only when the thermal inertia is sufficiently high globally.
@@
This behavior can be qualitatively understood using Eqs. (6) and (7). For an annual forcing frequency of s-1, the corresponding values of for Cases 1, 2, 3, and 4 are approximately 0.014, 0.045, 0.14, and 0.049 WmK-1, respectively. Among these, only Case 3 exceeds the atmospheric damping coefficient WmK-1, obtained as the annual mean, indicating that subsurface heat conduction becomes significant only when thermal inertia is sufficiently high. On seasonal timescales, subsurface conduction is generally less important than atmospheric damping primarily due to Titan’s long orbital period (29.5 Earth years), which yields a small . This stands in contrast to the diurnal regime, wherein subsurface heat conduction can dominate the surface temperature variability at intermediate values of thermal inertia.
@@
Figure 3 presents the seasonal evolution of surface energy balance components and surface temperature near the Huygens landing site for the four simulation cases. Across all cases, the net SW flux varies seasonally from approximately 0.20 to 0.37 Wm-2, increasing from winter through summer. The sensible heat flux exhibits a broadly similar seasonal pattern, ranging from 0.15 to 0.30 Wm-2. Lower-thermal-inertia cases exhibit larger fluctuation amplitudes, driven by greater surface temperature variability. The net IR flux remains relatively steady at approximately 0.07 Wm-2 across seasons and simulation cases. The resulting net energy flux is smaller in the lower thermal inertia cases, where a stronger sensible heat flux efficiently removes much of the incoming surface energy.
@@
Surface temperature variations across seasons remain modest—under 1 K—and exhibit broadly similar patterns across all simulation cases. However, the amplitude of these variations is more strongly damped for higher thermal inertia. Applying Eqs. (6) and (7) to the simulated seasonal net energy flux amplitudes of , , and Wm-2 for Cases 1, 2, and 3, respectively, yields theoretical temperature variations of , 0.57, and 0.51 K. Here, is derived from means calculated over the final 30 simulation years where 10-Titan-day moving averages are used to remove short-term variations. These estimates are roughly consistent with the simulated surface temperature variations of , , and K, though some discrepancies remain. These differences likely stem from the assumption of a constant across seasons, whereas in reality the horizontal wind speed exhibits seasonal variability in the simulations. Despite discrepancies between and due to the simplification, Eqs. (6) and (7) effectively capture the dominant role of atmospheric damping in regulating surface temperature variations at seasonal timescales. For instance, omitting the atmospheric damping term leads to theoretical predictions of , 1.22, and 0.70 K, which significantly overestimate the temperature response compared to the simulations. For Case 4, the Wm-2 yields a theoretical surface temperature variation of K, nearly identical to that of Case 2.
@@ 4.1 Diurnal Variations
Figure 4 shows the diurnal evolution of PBL height across the four simulation cases near the Huygens landing site. In each case, the left panels present snapshots of potential temperature profiles at 9:00 AM, 12:00 PM, and 3:00 PM local solar time (blue), as well as 12:00 AM (purple), compared with in-situ Huygens probe measurements (Fulchignoni et al., 2005) taken around 10:00 AM (black). Surface heating and the deepening of the adiabatic layer due to dry convection during the day are evident in all cases. At 12:00 AM, a stable near-surface inversion forms in Case 1, consistent with the negative sensible-heat flux in Figure 1. As thermal inertia increases across cases, the 12:00 AM profiles become progressively less stable, consistent with reduced nocturnal cooling and a warmer surface. Notably, in Case 3 the 12:00 AM profile is similar to those at other local times, indicating that a near-surface adiabatic layer persists throughout the diurnal cycle under high thermal inertia. The middle panels show the corresponding temperature diffusion tendencies predicted by the model’s PBL parameterization, with arrows marking the profiles at 9:00 AM, 12:00 PM, and 3:00 PM corresponding to the inflection points in the left panels. The diffusion tendencies are generally positive, reflecting upward heat transfer by convection, but contain interspersed negative layers that create a noisy appearance. The evolution of these tendencies closely tracks the changes in the potential temperature profiles.
The right panels display the mean composite evolution of PBL depth, averaged over 30 Titan days during midsummer (–). We diagnose PBL height as the height above ground where the vertical eddy diffusivity first falls below . This diffusivity-threshold criterion is the default in TAM when the Mellor–Yamada scheme is used and it closely delineates the top of the near-surface adiabatic layer. The time series is smoothed with a Savitzky–Golay filter (window = 5 bins, polynomial order = 2) to suppress bin-to-bin noise while preserving diurnal phase and amplitude. Results are insensitive to reasonable choices of window (3–7 bins). These diagnosed PBL heights reflect variations in surface heating and the resulting convective mixing, as corroborated by the inflection point of the potential temperature profile and the temperature diffusion tendencies in the left and middle panels. Although the individual PBL-depth time series show variability with occasional outliers (e.g., Case 2), these are attributable to natural variability; averaging effectively suppresses them.
For the lowest thermal inertia surface (Case 1), the PBL grows from m to m by mid-afternoon, where uncertainty ranges correspond to one standard deviation from the mean. With increasing thermal inertia, baseline nighttime PBL depths rise to m and m, and maximum daytime depths reach m and m for Case 2 and Case 3, respectively. The significant increase in nighttime PBL depths for higher thermal inertia cases reflects reduced nighttime surface cooling, which allows the PBL to remain deeper (Figure 1). For Case 4, the PBL grows from m to m by mid-afternoon, closely matching Case 2, which employs a comparable thermal inertia. Daytime maximum PBL depths are deeper for higher-thermal-inertia surfaces, despite the smaller daytime sensible heat fluxes over these surfaces. This likely reflects that daytime depth is controlled by the integrated heat accumulation over the day rather than by instantaneous fluxes. Indeed, the PBL fluctuation amplitude—reflecting the magnitude of PBL growth—is greatest for the lowest thermal inertia case, consistent with its larger sensible heat flux variability. Differences in sensible heat flux lead to noticeable variations near the tropopause (100 mbar, 50 km altitude): Case 1 shows a higher temperature (71 K) than Case 3 (70 K), with Case 2 and 4 intermediate (not shown). This reflects the cumulative effect of sensible heat flux over time: low thermal inertia surfaces transfer more energy upward, warming the free atmosphere and raising tropopause temperature, whereas high thermal inertia surfaces limit this warming.
@@ 4.2 Seasonal Variations
Figure 5 presents the seasonal evolution of the PBL height across the four simulation cases. The left panels in each case display global zonal means compiled over 10 Titan years, while the right panels show the 10-Titan-year mean composite PBL depth near the Huygens landing site, where Savitzky–Golay smoothing (30 Titan-day window, third order) is applied to emphasize the seasonal amplitude. The global seasonal structure of the PBL closely mirrors the surface temperature patterns shown in Figure 2, indicating that PBL depth variations are primarily governed by the seasonal migration of regions with peak surface heating. The hemispheric asymmetry in the PBL structure is evident, as it is in the surface temperature, due to orbital eccentricity.
In Case 1, the diagnosed PBL height at the Huygens landing site ranges from approximately 180 m during southern winter to about 1,100 m between the southern vernal () and autumnal () equinoxes. In Case 2, it ranges from 200 m to 1,320 m, while in Case 3, it extends from 290 m to 2,300 m. The increase in baseline (minimum) PBL depths for higher thermal inertia cases is modest, whereas the increase in maximum depths is more pronounced. In both cases, the elevated PBL heights for higher thermal inertia simulations are consistent with the diurnal-timescale behavior: a reduced contrast between minimum and maximum surface temperatures produces a more steady temperature structure that allows greater energy buildup and formation of a deeper PBL.
@@
Figure 6 presents global seasonal mean maps of the diagnosed PBL depth for four distinct seasons, using Case 4 to illustrate how local variations in thermal inertia influence PBL depth. The maximum depth is around 1,800 m, and structural differences across longitudes are apparent due to spatial variations in thermal inertia. For instance, an elevated PBL depth is observed in the Xanadu region (approximately 250°E near the equator) across the seasons, where the assumed thermal inertia corresponds to hummocky terrain (1,962 TIU), compared to dune fields (236 TIU) and plains (601 TIU) at similar latitudes elsewhere. The imprint of this enhancement is especially pronounced during the southern summer and winter solstices. Additionally, elevated PBL depths are evident in polar regions—near 90°S, 180°E during southern summer and near 70°N, 40°E during southern winter, attributable to the presence of “convective lakes” with extremely high thermal inertia (18,090 TIU) (MacKenzie et al., 2019). These results suggest that locally varying thermal inertia strongly influences the seasonal evolution of the PBL depth.
@@
Figure 7 shows the zonal and meridional wind profiles at the model grid point closest to the Huygens landing site for the four simulation cases, compared with the Huygens probe observations (Bird et al., 2005; Karkoschka, 2016). Positive values in the zonal and meridional directions correspond to eastward and northward motions, respectively. The seasonal mean is computed as a 30 Titan-day average centered around midsummer. During the period, the simulated wind profiles exhibit minimal diurnal variation (not shown). Across all cases, the simulations reproduce the observed wind structure in the lowest few kilometers, successfully capturing the key features of both the zonal and meridional components. The simulated near-surface wind magnitudes are on the order of – ms-1, consistent with results from previous Titan model intercomparison studies (Lora et al., 2019). While the simulated near-surface winds are centered near zero on average, the model does occasionally reproduce eastward winds consistent with observations; however, the strongest observed eastward values lie outside the model variability. This residual bias may reflect, at least in part, the omission of convective vertical momentum transport, though it could also indicate that the Huygens observation was atypical. As shown in Charnay et al. (2015), equatorial tropical methane storms during equinox can transport eastward winds and their associated momentum from higher altitudes downward, potentially contributing to the eastward orientation of the dune fields (Lorenz et al., 2006; Lorenz & Radebaugh, 2009).
Titan’s middle atmosphere exhibits eastward superrotation driven by upward and equatorward angular momentum transport, while friction keeps near-surface flow close to solid-body rotation; the resulting low-latitude momentum budget favors weak mean westward winds near the surface, which our simulations reproduce and which agree with prior Titan GCM studies (Flasar & Achterberg, 2009; Lewis et al., 2023; Lombardo & Lora, 2023; Lora et al., 2019). Above 4 km in the troposphere, however, our simulated zonal winds are systematically weaker than observed. This bias likely reflects differences in the simulated thermal structure: by thermal-wind balance, understated zonal temperature gradients yield weaker shear. In our dry runs, the zonal surface-temperature gradients peak toward the poles (Fig. 2), whereas including moist processes leads to more realistic gradients and winds (Lora et al., 2019). Although near-surface differences among Cases 1–4 are small, the stratospheric response varies substantially (not shown): peak speeds reach in Cases 1, 2, and 4 but only in Case 3, suggesting that global surface-temperature structure—set here by thermal inertia—can imprint on the circulation aloft. A full analysis of the vertical momentum pathways responsible for this sensitivity is left for future work, while our focus here is the lowest few kilometers where boundary-layer processes dominate and model–data agreement is generally good.
@@
For meridional winds, the flow near the surface is predominantly southward, transitioning to northward at altitudes around 2 km. This vertical wind reversal suggests the presence of a cross-equatorial Hadley circulation, characterized by rising motion in the southern summer hemisphere and descending motion in the northern winter hemisphere. As a result, there is northward return flow aloft and a compensating southward return flow near the surface at the latitude of the Huygens landing site (Lora et al., 2025). Wind magnitudes—both zonal and meridional—are generally smaller in the higher thermal inertia cases, although the differences are minimal. This likely reflects weaker temperature gradients, as higher thermal inertia surfaces dampen surface temperature contrasts (Figure 2), thereby reducing the thermal forcing that drives the atmospheric circulation.
@@ 5 PBL Characteristics at the Dragonfly Landing Site
In this section, we present the same analysis of surface temperature and PBL structure as in the previous sections, but for the planned Dragonfly landing site near Selk Crater (E, N) during southern midsummer () (Barnes et al., 2021; Lorenz et al., 2021). For Case 4, the dunes terrain type, with a thermal inertia of TIU, is assumed at the surface (MacKenzie et al., 2019).
Figure 8 shows the diurnal energy balance components at the Dragonfly landing site. Maximum insolation is reduced compared to the Huygens landing site, as the Dragonfly site experiences northern winter conditions while the Huygens site corresponds to southern summer. This results in a lower mean surface temperature at the Dragonfly site, reduced by approximately 0.1 K relative to the Huygens site. Nevertheless, the amplitude of diurnal surface temperature variations remains comparable to that at the Huygens site. For Case 4, the simulated diurnal amplitude is K.
Figure 9 shows the diurnal PBL structures at the Dragonfly landing site. Due to the reduced surface energy input and lower temperatures during this winter season—compared to the midsummer conditions at the Huygens landing site—the diagnosed PBL is shallower, with maxima of m, m, m, and m across the four cases, in contrast to 1,100-1,400 m at the Huygens site. At night, the baseline minimum PBL depths are m, m, m, and m for Cases 1–4, respectively. This results in much-reduced diurnal amplitudes (minimum to maximum) of 300–400 m, compared with 800–1,100 m at the Huygens site.
Figure 10 shows the seasonal evolution of PBL depth at the Dragonfly landing site. As at the Huygens landing site, higher thermal inertia surfaces produce deeper PBLs, with maxima of , , , and m and minima of , , , and m for Cases 1–4, respectively. However, a key difference is observed in the seasonal symmetry. At the Dragonfly site, the PBL depth exhibits a more symmetric seasonal pattern, owing to its near-equatorial location compared to that of Huygens: maxima occur near the equinoxes ( and ), and minima near the solstices ( and ). In contrast, the Huygens site displays an asymmetric pattern: a pronounced minimum at the southern winter solstice (), with a higher but still suppressed depth at the southern summer solstice (), relative to the equinoctial maxima. This asymmetry is attributed to Saturn’s orbital eccentricity, as illustrated in Figure 2 and the left panels of Figure 5.